Chris, age 68, has an annuity upon which he will collect $8,000 in 2018. He originally invested $100,000 in the annuity. The annuity plan was set up for him to collect $40,000 annually for 10 years. How much of his annuity income is taxable in 2018? Use the general rule method to calculate the taxable amount. 1. 10,000 2. 6,000 3. 30,000 4. 2,000 5. 8,000
Answer #1
Total amount to be received 400000 =40000*10
Total amount taxable 300000 =400000-100000
Percentage of annuity taxable 75% =300000/400000
Taxable amount in 2018 6000 =8000*75%
Option 2 6,000 is correct